If the perimeter of a rectangle is to be no greater than 300 in.,and the length must 125 in. What is the maximum width of the rectangle?
Answer by checkley71(8403)

300=2L+2W OR 300=2*125+2W OR 300=250+2W OR 50=2W OR W=50/2 OR W=25